OpenStudy (anonymous):

Please help solve: radical 40 minus 2 radical 5/radical 10

OpenStudy (anonymous):

whats \[2\sqrt{10}\times \sqrt{10}\]

OpenStudy (anonymous):

help please...anyone! im really confused:/

OpenStudy (anonymous):

20

OpenStudy (anonymous):

wait so how do you solve the whole problem and what would the answer be?

OpenStudy (anonymous):

2 rad 10 x rad 10 = 10 x 2 = 20

OpenStudy (anonymous):

ok go on(:

OpenStudy (anonymous):

help me finish the whole problem

OpenStudy (anonymous):

oh ithought its a dffrnt problem

OpenStudy (anonymous):

no its the problem from the original post(:

OpenStudy (anonymous):